OpenStudy (anonymous):

How do I simplify 4c^2(5c^4-3)

OpenStudy (luigi0210):

Distribute it out

OpenStudy (anonymous):

How do I distribute it?

OpenStudy (luigi0210):

\[4c^2(5c^4-3)=(4c^3)(5c^4)-(4c^3)(3)\]

OpenStudy (luigi0210):

*2

OpenStudy (luigi0210):

\[a^x(a^y)=a^{x+y}\] \[(a^x)^y=a^{x*y}\]

OpenStudy (anonymous):

Alright. So, when I multiply (4c^2) (5c^4) the power because 6? And do I just multiply 4 by 3 or all of it? Sorry, I just got back to school and I'm lost.

OpenStudy (anonymous):

becomes*

OpenStudy (luigi0210):

Yes, you are correct

OpenStudy (anonymous):

Is it 20c^6-12c^2?

OpenStudy (luigi0210):

Yup, that's all you have to do

OpenStudy (anonymous):

Sweet! Thanks

OpenStudy (luigi0210):

Yup!

OpenStudy (anonymous):

Wait. Do you have to subtract the two after that or is it done?

OpenStudy (luigi0210):

Well the equation you have is like: x^3-x, tell me, can you subtract those?

OpenStudy (anonymous):

I don't think so because the power makes it different than the other x so we don't know what it is to subtract, right?

OpenStudy (luigi0210):

yup, so you have to leave it as it is.